OpenStudy (anonymous):

1/2x+1/8x=3/4x-5 please somebody

OpenStudy (anonymous):

I assume you're solving for x, so first make all the fractions alike 4/8x+1/8x= 6/8x-5 --- now combine the two like terms to get 5/8x=6/8x -5 --- now move the 6/8x across the sign -6/8x -6/8x --- -1/8x = -5 --- solve for x by multiplying both sides by -8 x=40